# supplement

### v1

@@ D.6 Proofs of the coverage rate guarantees
In this section, we prove Proposition 4.1 and Theorem 5.2. First, we present a general theory that extends these results to a general design of censoring times. This general result builds on the proof of [15, Theorem 3].
Theorem D.1 (General validity).
@@
$w_{\tau}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)={\mathbb{P}}\bigl{[}\hat{q}_{\tau}(X_{i})\leq C_{i}\bigm{|}\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr{]},$ (12)
and that there exists a constant ${\gamma}_{\tau}>0$ such that the weights satisfy ${w}_{\tau}(x)\leq{\gamma}_{\tau}$ for $P_{X}$-almost all $x$. Consider the following estimated miscoverage rate:
$\hat{\alpha}(\tau)=\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}w_{\tau}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\;\mathbb{I}\bigl{\{}\hat{q}_{\tau}(X_{i})\leq C_{i}\bigr{\}}\;\mathbb{I}\bigl{\{}T_{i}<\hat{q}_{\tau}(X_{i})\bigr{\}}$ (13)
@@
Above, the search space $\mathcal{T}$ is formulated as in Section 4, or using $\mathcal{T}\cap[0,{\tau_{\textup{prior}}}]$, as in the proposed Adaptive method from Section 5. We remark that for $\hat{\tau}$ to be well defined, we assume there exists $\tau^{\prime}\in\mathcal{T}$ such that $\hat{\alpha}(\tau^{\prime})\leq\alpha$ This assumption can be trivially satisfied by setting $\hat{q}_{0}(X_{i})=0$. Then, with probability at least $1-\delta$ over the draws of $\mathcal{D}$, the LPB $\hat{L}(x)=\hat{q}_{\hat{\tau}}(x)$ satisfies
@@
$\begin{split}&\mathbb{P}(\hat{\alpha}(\tau(\alpha+\Delta)+\varepsilon)\leq\alpha\mid{\mathcal{I}_{1}})\\ =&\mathbb{P}(\hat{\alpha}(\lambda)\leq\alpha\mid{\mathcal{I}_{1}})\\ =&\mathbb{P}\left(\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}w_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\;\mathbb{I}\bigl{\{}\hat{q}_{\lambda}(X_{i})\leq C_{i}\bigr{\}}\;\mathbb{I}\bigl{\{}T_{i}<\hat{q}_{\lambda}(X_{i})\bigr{\}}\leq\alpha\bigg{|}{\mathcal{I}_{1}}\bigg{.}\right)\\ =&\mathbb{P}\left(\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}w_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\;\mathbb{I}\bigl{\{}\hat{q}_{\lambda}(X_{i})\leq C_{i}\bigr{\}}\;\mathbb{I}\bigl{\{}T_{i}<\hat{q}_{\lambda}(X_{i})\bigr{\}}-\alpha\leq 0\bigg{|}{\mathcal{I}_{1}}\bigg{.}\right)\end{split}$ (19)
@@
(19)$&\leq\mathbb{E}\left(\exp\left(\alpha-\frac{t}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}w_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\;\mathbb{I}\bigl{\{}\hat{q}_{\lambda}(X_{i})\leq C_{i}\bigr{\}}\;\mathbb{I}\bigl{\{}T_{i}<\hat{q}_{\lambda}(X_{i})\bigr{\}}\right)\bigg{|}{\mathcal{I}_{1}}\bigg{.}\right)\end{split}$ (20)
@@
(20)$&\leq\exp\left(\frac{\gamma^{2}_{\lambda}t^{2}}{8|{\mathcal{I}_{2}}|}\right)\cdot\mathbb{E}\left(\exp\left(\alpha-\frac{t}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}w_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\;w_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)^{-1}\;\mathbb{I}\bigl{\{}T_{i}<\hat{q}_{\lambda}(X_{i})\bigr{\}}\right)\bigg{|}{\mathcal{I}_{1}}\bigg{.}\right)\\ &=\exp\left(\frac{\gamma^{2}_{\lambda}t^{2}}{8|{\mathcal{I}_{2}}|}\right)\cdot\mathbb{E}\left(\exp\left(\alpha-\frac{t}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}\mathbb{I}\bigl{\{}T_{i}<\hat{q}_{\lambda}(X_{i})\bigr{\}}\right)\bigg{|}{\mathcal{I}_{1}}\bigg{.}\right)\end{split}$ (22)
@@
(22)$\leq\exp\left(\frac{(\gamma^{2}_{\lambda}+1)t^{2}}{8|{\mathcal{I}_{2}}|}\right)\mathbb{E}\left(\exp\left(\frac{t}{|{\mathcal{I}_{2}}|}\cdot\sum_{i\in{\mathcal{I}_{2}}}(\alpha-p_{\lambda}(X_{i}))\right)\bigg{|}{\mathcal{I}_{1}}\bigg{.}\right)\end{split}$ (24)
@@
(25)$&=\mathbb{E}\left(\exp\left(\frac{2t}{|{\mathcal{I}_{2}}|}\cdot\sum_{i\in{\mathcal{I}_{2}}}(\alpha-p_{\lambda}(X_{i}))\right)\bigg{|}{\mathcal{I}_{1}}\bigg{.}\right)\\ \leq&\mathbb{E}\left(\exp\left(\frac{2t}{|{\mathcal{I}_{2}}|}\cdot\sum_{i\in{\mathcal{I}_{2}}}(\mathbb{P}(T<\hat{q}_{\lambda}(X)\mid{\mathcal{I}_{1}})-\Delta-p_{\lambda}(X_{i}))\right)\bigg{|}{\mathcal{I}_{1}}\bigg{.}\right)\\ =&\exp(-2t\Delta)\mathbb{E}\left(\exp\left(\frac{2t}{|{\mathcal{I}_{2}}|}\cdot\sum_{i\in{\mathcal{I}_{2}}}(\mathbb{P}(T<\hat{q}_{\lambda}(X)\mid{\mathcal{I}_{1}})-p_{\lambda}(X_{i}))\right)\bigg{|}{\mathcal{I}_{1}}\bigg{.}\right)\\ \leq&\exp\left(\frac{t^{2}}{2|{\mathcal{I}_{2}}|}-2t\Delta\right).\end{split}$ (27)
@@
(19)$\leq&\exp\left(\frac{t^{2}}{2|{\mathcal{I}_{2}}|}-2t\Delta+\frac{(\gamma^{2}_{\lambda}+1)t^{2}}{8|{\mathcal{I}_{2}}|}+\frac{\gamma^{2}_{\lambda}t^{2}}{8|{\mathcal{I}_{2}}|}\right)\\ =&\exp\left(\frac{(2\gamma^{2}_{\lambda}+5)t^{2}}{8|{\mathcal{I}_{2}}|}-2t\Delta\right)\end{split}$ (28)
@@
(24)$\leq&\exp\left(\frac{(2\gamma^{2}_{\lambda}+5)(\frac{(8+2\sqrt{14})|{\mathcal{I}_{2}}|\Delta}{2\gamma^{2}_{\lambda}+5})^{2}}{8|{\mathcal{I}_{2}}|}-\frac{(16+4\sqrt{14})|{\mathcal{I}_{2}}|\Delta}{2\gamma^{2}_{\lambda}+5}\Delta\right)\\ =&\exp\left(\frac{(15+4\sqrt{14})|{\mathcal{I}_{2}}|\Delta^{2}}{(2\gamma^{2}_{\lambda}+5)}-\frac{(16+4\sqrt{14})}{2\gamma^{2}_{\lambda}+5}|{\mathcal{I}_{2}}|\Delta^{2}\right)\\ =&\exp\left(-\frac{1|{\mathcal{I}_{2}}|\Delta^{2}}{2\gamma^{2}_{\lambda}+5}\right)\\ \leq&\exp\left(-\log\left(\frac{1}{\delta}\right)\right)\\ =&\delta\\ \end{split}$ (30)
@@
Since the proposed methods construct censoring times that satisfy the conditional independence assumption of Theorem D.1, the proofs of Proposition 4.1 and Theorem 5.2 follow immediately from this theorem with the corresponding definitions of the censoring times and the weights.
@@ E.1.1 Additional experiments
We repeated the experiment for the following values of average sampling budget per prompt
$B/|I_{2}|:{10,25,50,100,200,300,600,1200}.$
In all experiments, we used the same prior quantile $\tau_{\mathrm{prior}}=10^{-1/4}$, with trimming threshold $M$ set so that $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right)=10$, as in Section 6.1.
Figure 3 presents the performance of the methods introduced in this work for 20 data splits. Across every budget level, the performances closely match those from our original fixed-split experiment from Figure 1 in Section 6.1.
@@
Figure 4: Results of synthetic experiments as a function of average budget per prompt $B/|{\mathcal{I}_{2}}|$. Top: Coverage (target $100\cdot(1-\tau)$%; gray dashed line). Bottom: Mean LPB. Shaded regions represent the standard deviation over 20 runs.
Finally, we turn to study the effect of the maximal weight size $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right)$ on the stability and power of the Trimmed and Optimized calibration methods. Figure 5 compares the two methods under a fixed per-prompt budget of ${B}/{|{\mathcal{I}_{2}}|}=1000$. For each weight size, we set the threshold $M={B\gamma}/{|{\mathcal{I}_{2}}|}$. As can be seen in this figure, the size of the LPBs we construct increases with $\gamma$. However, this comes at the cost of increasing the coverage variance. Observe how the variance of Optimized method for $\gamma=100$ is similar to that of Trimmed method for $\gamma=10$, but the former yields much more informative LPBs under these corresponding $\gamma$ values. This is a result of the variance-reducing optimization objective in (8).
@@ E.2 Real-Data Experiments
Efficient prompt-parallel sampling
To efficiently generate outputs during our calibration procedures, we implement a prompt-parallel generation algorithm on top of the vLLM package [33]. In each iteration, we gather a batch of $1500$ prompts ${X_{i}}$ along with their corresponding censoring budgets ${C_{i}}$, and distribute them across multiple GPUs. The pipeline iteratively performs the following steps in parallel for a batch of prompts, until all prompts in the calibration set are processed:
@@
We employ the ModernBERT-base [61] model. We use Adam optimizer [29] to minimize the loss function in (33), with a learning rate of 3e-5, and a batch size of $1500$. We train the model for 10 epochs and use early stopping based on the validation set.
@@ E.3.1 Implementation details of the Naive calibration procedure
E.4 Machine specifications
The hardware and OS used for the experiments are as follows.
  • CPU: AMD EPYC 7443 24-Core Processor
  • GPU: NVIDIA RTX A6000
  • OS: Ubuntu 20.04

### v4

@@ D.6 Proofs of the coverage rate guarantees
In this section, we present and prove the formal version Theorem 4.1 which builds on the proof of (Gui et al., 2024, Theorem 3).
Theorem D.1 (General validity, formal).
@@
$w_{\tau}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)=1/{\mathbb{P}}\bigl[\hat{q}_{\tau}(X_{i})\leq C_{i}\bigm|\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr].$ (13)
Above, $\hat{q}_{\tau}(X_{i})$ can be either the estimated quantile of $T_{i}\mid X=X_{i}$ presented in Section 3, or its trimmed version $\hat{f}_{\tau}(X_{i})$ from Section 5.3. We assume that there exists a constant ${\gamma}_{\tau}>0$ such that the weights satisfy ${w}_{\tau}(x)\leq{\gamma}_{\tau}$ for $P_{X}$-almost all $x$. Consider the following estimated miscoverage rate:
$\hat{\alpha}(\tau)=\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}w_{\tau}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\;\mathbb{I}\bigl\{\tilde{T}_{i}<\hat{q}_{\tau}(X_{i})\leq C_{i}\bigr\}$ (14)
@@
Above, the search space $\mathcal{T}$ is formulated as in Section 4, or using $\mathcal{T}\cap[0,{\tau_{\textup{prior}}}]$, as in the proposed Adaptive method from Section 5.2. We remark that for $\hat{\tau}$ to be well defined, we assume there exists $\tau^{\prime}\in\mathcal{T}$ such that $\hat{\alpha}(\tau^{\prime})\leq\alpha$ . This assumption can be trivially satisfied by setting $\hat{q}_{0}(X_{i})=0$. Then, with probability at least $1-\delta$ over the draws of $\mathcal{D}$, the LPB $\hat{L}(x)=\hat{q}_{\hat{\tau}}(x)$ satisfies
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Since the proposed methods construct censoring times that satisfy the conditional independence assumption of Theorem D.1, their validity follow immediately from this theorem with the corresponding definitions of the censoring times and the weights.
@@ E.1 Synthetic experiments
Calibration method parameters
In all experiments, unless stated otherwise, we used the prior quantile $\tau_{\mathrm{prior}}=10^{-1/4}$, with trimming threshold $M$ set so that $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right)=10$.
@@ E.1.1 Additional experiments
We repeated the experiment for the following values of average sampling budget per prompt $B/|I_{2}|$: 10,25,50,100,200,300,600,1200{10,25,50,100,200,300,600,1200}. In all experiments, we used the same prior quantile $\tau_{\mathrm{prior}}=10^{-1/4}$, with trimming threshold $M$ set so that $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right)=10$, as in Section 6.1.
Figure 3 presents the performance of the methods introduced in this work for 20 data splits. Across every budget level, the performances closely match those from our original fixed‐split experiment from Figure 1 in Section 6.1.
@@
Figure 4: Results of synthetic experiments as a function of average budget per prompt $B/|{\mathcal{I}_{2}}|$. Top: Coverage (target level indicated by a gray dashed line). Bottom: Mean LPB. Shaded regions represent the standard deviation over 20 runs.
Finally, we turn to study the effect of the maximal weight size $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right)$ on the stability and power of the Trimmed and Optimized calibration methods. Figure 5 compares the two methods under a fixed per‐prompt budget of ${B}/{|{\mathcal{I}_{2}}|}=1000$. For each weight size, we set the threshold $M={B\gamma}/{|{\mathcal{I}_{2}}|}$. As can be seen in this figure, the size of the LPBs we construct increases with $\gamma$. However, this comes at the cost of increasing the coverage variance. Observe how the variance of Optimized method for $\gamma=100$ is similar to that of Trimmed method for $\gamma=10$, but the former yields much more informative LPBs under these corresponding $\gamma$ values. This is a result of the variance-reducing optimization objective in (10).
@@
We compare the proposed Optimized method with an oracle quantile regression model on our synthetic dataset. The oracle model outputs the true conditional quantile of $T\mid X$ as the LPB, and therefore achieves the target coverage rate of $1-\alpha=90\%$ by definition. These oracle LPBs are also the most informative possible by construction. Figure 6 reports the performance of the Uncalibrated quantile regression model, Naive, Basic, Trimmed, Optimized, and the oracle quantile regression model. As expected, the oracle model achieves the nominal coverage level. Notably, our Optimized method produces LPBs of comparable size to the oracle as the budget increases, demonstrating its efficiency in approaching the best achievable performance even without access to the oracle quantiles.
Figure 6: Synthetic data experiment with an oracle quantile regression model. Top: Coverage (target $90\%$%; gray dashed line). Bottom: Mean LPB. Shaded regions represent the standard deviation over 20 runs.
We compare our approach with a calibration method that has an infinite budget on our synthetic data. With an unlimited budget, we can observe an unsafe event for every calibration sample and thus can employ classic conformal prediction (Vovk et al., 2005, Papadopoulos, 2008) without data re-weighting to obtain LPBs that achieve $1-\alpha=90\%$ coverage rate. Specifically, we estimate the miscoverage rate similarly to (6), but here, we use all calibration points:
$\hat{\alpha}^{\text{ib}}(\tau)=\frac{1}{|{\mathcal{I}_{2}}|+1}\sum_{i\in\mathcal{I}_{2}}\mathbb{I}\bigl\{T_{i}<\hat{q}_{\tau}(X_{i})\bigr\}$ (34)
and construct the LPB as in (7) with the above $\hat{\alpha}^{\text{ib}}$.
We summarize the performance of all methods in Figure 7. The results reveal that the infinite budget method achieves the desired coverage rate with lower variability than the alternatives. Moreover, our Optimized approach produces LPBs comparable in size to those of the infinite-budget calibration as the budget increases, demonstrating its effectiveness in approaching optimal performance even under budget constraints.
Figure 7: Synthetic data experiment with a calibration method that has an infinite budget. Top: Coverage (target 90%; gray dashed line). Bottom: Mean LPB. Shaded regions represent the standard deviation over 20 runs.
@@ E.2 Real-Data Experiments
Calibration method parameters
In all experiments, unless stated otherwise, we used the prior quantile $\tau_{\mathrm{prior}}=10^{-1/4}$, with trimming threshold $M$ set so that $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right)=2$.
Efficient prompt‐parallel sampling
To efficiently generate outputs during our calibration procedures, we implement a prompt-parallel generation algorithm on top of the vLLM package (Kwon et al., 2023). In each iteration, we gather a batch of $1500$ prompts $X_{i}$} along with their corresponding censoring budgets $C_{i}$}, and distribute them across multiple GPUs. The pipeline iteratively performs the following steps in parallel for a batch of prompts, until all prompts in the calibration set are processed:
@@
We employ the ModernBERT-base (Warner et al., 2024) model. We use Adam optimizer (Kingma & Ba, 2015) to minimize the loss function in (35), with a learning rate of 3e-5, and a batch size of $1500$. We train the model for 10 epochs and use early stopping based on the validation set.
@@ E.3.1 Implementation details of the Naive calibration procedure
E.4 Machine’s spec
The computational infrastructure used in the experiments are:
  • CPU: Intel(R) Xeon(R) CPU E5-2683 v4 @ 2.10GHz, Intel(R) Xeon(R) Gold 5318Y CPU @ 2.10GHz, Intel(R) Xeon(R) Gold 6336Y CPU @ 2.40GHz.
  • GPU: NVIDIA A40, NVIDIA TITAN X (Pascal), NVIDIA 2080 TI, NVIDIA RTX 2060 SUPER.
  • OS: Ubuntu 20.04.6.